OpenStudy (anonymous):

Solve the following linear expressions by elimination: 2x+y+z=1 x-y+2z=5 3x+4y-z=0

OpenStudy (anonymous):

may you help me out wtih this?

OpenStudy (jdoe0001):

have you done elimination before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get 21x+21z=42 -21x-21z=-60 0 does not equal -18

OpenStudy (anonymous):

so is the solution trivial?

OpenStudy (jdoe0001):

ok.... so..... lemme show you how to do it for 3 variables so we take 2 functions at a time eliminate one of the 3 variables then we end up with a 2 variable set then we do the same thing again then we end up with 2 variables system of equations 2x+y+z=1 <---- let's use this one x-y+2z=5 <---- and this one this time 3x+4y-z=0

OpenStudy (jdoe0001):

gimme a sec

OpenStudy (anonymous):

ty, jdoe, you are the best

OpenStudy (anonymous):

you still there?

OpenStudy (jdoe0001):

yes hehee, I was typing..... ...just not here :)

OpenStudy (anonymous):

ty

OpenStudy (jdoe0001):

\(\begin{array}{llrll} 2x+y+z=1& \times -2\implies &-4x-2y-2z=-2\\ x-y+2z=5&&x\quad -y\quad +2z=5\\ \hline\\ &&-3x\quad -3y\quad +0z=3\\ \quad \\ \bf \color{blue}{-3x-3y=3} \end{array}\)

OpenStudy (anonymous):

hmm

OpenStudy (jdoe0001):

so that's our first 2 variable function now for our next one 2x+y+z=1 <-- let's use this one again x-y+2z=5 3x+4y-z=0 <--- and this one this time one sec

OpenStudy (anonymous):

same problem, it cancels out to 0 does not equal 18 :/

OpenStudy (jdoe0001):

\(\begin{array}{rllrll} 2x+y\quad +z=1&\textit{notice we don't have to multiply by anything}\\ 3x+4y-z=0\\ \hline\\ 5x+5y+0z=1\\ \quad \\ \bf \color{blue}{5x+5y=1} \end{array}\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but

OpenStudy (anonymous):

then you have to add up -3x-3y=3 +5x+5y=1

OpenStudy (anonymous):

and they cancel out to 0 does not equal 18

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

lemme check that

OpenStudy (anonymous):

ty, jdoe, you are a HERO

OpenStudy (jdoe0001):

hmm you're right.... it ends up as 0 = 18 that means that the system is inconsisten, and thus it has no solution, meaning the 3 lines do not touch each other

OpenStudy (anonymous):

jdoe

OpenStudy (anonymous):

if it has one solution only, is it Independent and consistent?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ty

OpenStudy (jdoe0001):

yw